OpenStudy (anonymous):

subtract, simplify if possible 17/e-1 - 17/1-e

OpenStudy (anonymous):

note that 1-e = -(e-1)

jhonyy9 (jhonyy9):

17/(e-1) = (-17)/(1-e) so than -17/(1-e) - 17/(1-e) =

jhonyy9 (jhonyy9):

???

jhonyy9 (jhonyy9):

hope that you will continue

jhonyy9 (jhonyy9):

when there is the same denominators

jhonyy9 (jhonyy9):

you need just assum the numerators yes ???

OpenStudy (anonymous):

sorry i was away for a min

jhonyy9 (jhonyy9):

so than will be (-17-17)/(1-e) = -34/(1-e) yes ?

jhonyy9 (jhonyy9):

right ???

OpenStudy (anonymous):

i have no idea.. i dont know how to o this stuff

jhonyy9 (jhonyy9):

there is denomintor (e-1) and (1-e) yes ?

OpenStudy (anonymous):

yes

jhonyy9 (jhonyy9):

but we know that two fraction can be assumed when there are denominator the same or we need get or calcule the denominator common

jhonyy9 (jhonyy9):

right ?

OpenStudy (anonymous):

yes

jhonyy9 (jhonyy9):

so than because from (e-1) we can make (1-e) if ,so you know how ?

jhonyy9 (jhonyy9):

comen i like to learn to you

OpenStudy (anonymous):

im confused.. sorry I dont understand this stuff or what im supposed to be doing..

jhonyy9 (jhonyy9):

so there is like denominators (e-1) and (1-e)

OpenStudy (anonymous):

yes

jhonyy9 (jhonyy9):

how can you assume two freactions ?

OpenStudy (anonymous):

no idea...lol

jhonyy9 (jhonyy9):

so looke please 1/2 + 2/3 = ?

OpenStudy (anonymous):

3/5

jhonyy9 (jhonyy9):

no

jhonyy9 (jhonyy9):

we need calcule the denominator common for 2 and 3

jhonyy9 (jhonyy9):

do you know how ?

OpenStudy (anonymous):

the common factor of 2 and 3?? that is 6 right?

jhonyy9 (jhonyy9):

yes right

jhonyy9 (jhonyy9):

so for 1-e and e-1

jhonyy9 (jhonyy9):

???

OpenStudy (anonymous):

they are the same to me?

OpenStudy (anonymous):

1e

jhonyy9 (jhonyy9):

why ? for you is indefferent if you get 1-3 dollars or 3-1 dollars ?

OpenStudy (anonymous):

1-3 dollars is -2 dollars... 3-1 dollars is 2 dollars..

jhonyy9 (jhonyy9):

yes so firstly you need give me 2 dollars but secondly i need give to you 2 dollars

jhonyy9 (jhonyy9):

right not are the same

OpenStudy (anonymous):

i dont know how to subtract a letter from a letter.. lol

jhonyy9 (jhonyy9):

not is necessary

jhonyy9 (jhonyy9):

because you can make from e-1 if you multiply this by (-1) so how many will be ?

OpenStudy (anonymous):

2e

OpenStudy (anonymous):

-2e

jhonyy9 (jhonyy9):

no multiply (e-1)(-1)=

OpenStudy (anonymous):

-1e..

jhonyy9 (jhonyy9):

comen jimmyrep help me make to understand this girl the exercise

jhonyy9 (jhonyy9):

no you need multiply every terms by (-1)

OpenStudy (anonymous):

-1e and 1

jhonyy9 (jhonyy9):

so than will be (-1)(e) a

jhonyy9 (jhonyy9):

yes but how can you write this mathematicaly

jhonyy9 (jhonyy9):

not like -e+1

OpenStudy (anonymous):

-1e, 1e?

jhonyy9 (jhonyy9):

so than will be -e+1

jhonyy9 (jhonyy9):

what is equal 1-e

OpenStudy (anonymous):

e

jhonyy9 (jhonyy9):

than if you see will be the same every two denominators

jhonyy9 (jhonyy9):

17/(-(1-e)) - 17/(1-e)

jhonyy9 (jhonyy9):

17/(-(1-e)) = (-17)/(1-e)

jhonyy9 (jhonyy9):

so than there is (-17)/(1-e) - 17/(1-e)

jhonyy9 (jhonyy9):

what are two fraction the same

jhonyy9 (jhonyy9):

so how many will be ?

jhonyy9 (jhonyy9):

-2/3 -2/3 = ???